OpenStudy (osuxbucks15):

How many grams of iron (3) carbonate contains 17g of carbonate ions?

OpenStudy (preetha):

I could answer this and help you, but your name is offensive to OS and its users.

OpenStudy (anonymous):

i don't understand how the name is offensive?

OpenStudy (osuxbucks15):

ok so iron 3 carbonate is Fe2(CO3)3

OpenStudy (osuxbucks15):

right?

OpenStudy (anonymous):

give me one second

OpenStudy (osuxbucks15):

alright

OpenStudy (preetha):

My Public Apology to OSU!

OpenStudy (anonymous):

do you mind helping @Preetha

OpenStudy (anonymous):

yes i do think you need molar mass i'm researching a bit on this just so you get a good answer

OpenStudy (osuxbucks15):

thank you

OpenStudy (preetha):

I think you are trying to calculate how many g of Fe2(CO3)3 would be needed to get 17 g of CO3. If so then here it is: For every one mole of the Iron carbonate, you have 3 moles of CO3. So for 17 g of CO3 you need: 17 g CO3 x 1 mole of CO3/molar mass of CO3 x 3 moles CO3/1 mole Fe2(CO3)3 x molar mass of Fe2(CO3)3 /1 mole of Fe2(CO3)3. Does that make sense?

OpenStudy (osuxbucks15):

ok so the molar mass would be...55.8(2) + 12(3) + 16(9)=291.6

OpenStudy (anonymous):

that sounds correct @Preetha :)

OpenStudy (osuxbucks15):

17 x 6.02x10^23 =1.02x10^25

OpenStudy (osuxbucks15):

then...1.02x10^25/ 180=5.7x10^22, is this right so far?

OpenStudy (preetha):

I dont know why you are multiplying by avogadro's number OSU. In this step, 17 g CO3 x 1 mole of CO3/molar mass of CO3 you have to convert the g to moles by dividing by the molar mass of CO3

OpenStudy (osuxbucks15):

so u dont use 6.02x10^23

OpenStudy (osuxbucks15):

?

OpenStudy (preetha):

No. How do you convert 17g to moles of CO3?

OpenStudy (anonymous):

Molar mass of CO3 is 60.00907 ± 0.00004 g/mol oops

OpenStudy (anonymous):

sorry my things wrong

OpenStudy (osuxbucks15):

17/60=.28

OpenStudy (anonymous):

yes use my second eqution mol X g/mol = grams grams / g/mol = moles.

OpenStudy (osuxbucks15):

do i multiply .28 by 3

OpenStudy (anonymous):

grams/ g/mol= moles to convert

OpenStudy (anonymous):

wait one sec

OpenStudy (anonymous):

preetha am i wrong?

OpenStudy (anonymous):

I'm sorry my helps not that good on this question I slightly forget learning this but i think you do multiply by three

OpenStudy (osuxbucks15):

ok so .28 x 3 = .84/291.6....

OpenStudy (osuxbucks15):

is this right?

OpenStudy (osuxbucks15):

so the answer is .003?

OpenStudy (anonymous):

yes I would asume so! I'd check with some peers too are you in a home school?

OpenStudy (osuxbucks15):

nope

OpenStudy (anonymous):

then please check with someone I really hope that you get this correct I'm sorry I couldn't help more

OpenStudy (osuxbucks15):

its fine thank you

OpenStudy (anonymous):

do you have anything else? I can try to help you

OpenStudy (osuxbucks15):

nope, thank you

OpenStudy (anonymous):

:) sorry bye bye